# Sparse identification of quasipotentials via a combined data-driven method

### v1

@@ 3.2 Example 2: Nanomechanical graphene resonator
Long-term prediction. Here we analyze the accuracy of the prediction for the identified dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$. Given $1000$ observed trajectories of the system (23) with initial states sampled from $\{\mathbf{x}\in\Omega:V_{\theta}(\mathbf{x})<\tau\}$, we generate trajectories by evolving the identified dynamics from the states. The statistics of the errors as in Eq. (18) for the generated trajectories are $0.0242\pm 0.0258$. Fig. 5 (a) shows a comparison of the generated trajectories with the observed ones, demonstrating the capacity to accurately capture the underlying dynamical evolution.
Invariant distribution. With the identified quasipotential $U^{\text{Symb}}(P,Q)$ of Eq. (24), we compute an explicit invariant distribution for the nanomechanical resonator modelled by  (23) when perturbed by a random noise. It is important to investigate how fluctuations affect the out-of-equilibrium dynamics of nonlinear resonators because atomically thin resonators are very sensitive to environmental noise. The invariant distribution of the randomly perturbed system is approximated by $p_{\epsilon}(P,Q)=Z_{\epsilon}^{-1}\exp(-U^{\text{Symb}}(P,Q)/\epsilon)$. Here the normalization constant is estimated by the integration over the domain $\Omega$. The evolution of the invariant distribution $p_{\epsilon}(P,Q)$ for different values of $\epsilon$, namely $10^{-9}$, $2\times 10^{-9}$ and $3\times 10^{-9}$, is showcased in Fig. 5 (b)-(d). While the predominance of the large oscillations of the nanomechanical resonator persists (motion around the large amplitude attractor - fixed point $A$), we observe a gradual smearing of the distribution in the phase space of the quadratures for larger $\epsilon$.
@@ 4 Conclusions
To conclude, we introduced a computational method for discovering explicit expressions of the quasipotential from observed data, which combines a neural network approximation and sparse regression. The efficacy of the method was illustrated using two model systems. The first example demonstrates the accuracy of the identified quasipotential, while the second example highlights the ability of the data-driven method to interpret nanomechanical vibrational systems of practical engineering interest.
A feature of the suggested approach is that it is directly applicable to both synthetic and experimental datasets since it only necessitates observing the nonlinear dynamics. The data-driven approach can be used to examine a wide variety of systems that undergo complex multistable behaviour, provided that sufficient data are available. Moreover, being based on neural networks and sparse regression makes the algorithm scalable to high-dimensional systems.
For the first time, explicit quasipotential expressions may be found without making any presumptions about the governing vector field. Because of this, we believe that the proposed approach is of interest to find a precise prediction of pertinent statistical values pertaining to the transition between metastable states. This new idea of determining a symbolic quasipotential function by coupling machine learning techniques offers a fully data-driven solution, and it demonstrates how recent developments in data science and machine learning open up intriguing opportunities in the realm of nonlinear dynamics.
Appendix A Initialization step for the problem of Eqs. (14) and (16)
@@ Appendix B Data sampling and neural networks training
Sampling data $\Omega$ $[-2,2]\!\times\![-1.5,1.5]\!\times\![-1.5,1.5]$ $[-0.02,0.02]\!\times\![-0.02,0.02]$
$N$ $5000$ $2000$
@@
$V_{\theta}^{\text{NN}}$ Structure $3$-$50$-$50$-$50$-$1$ $2$-$100$-$100$-$100$-$1$
@@
$\mathbf{g}_{\theta}^{\text{NN}}$ Structure $3$-$50$-$50$-$50$-$3$ $2$-$100$-$100$-$100$-$2$
@@
Training Optimizer Adam Adam
@@
Data points are collected from $N$ trajectories generated from numerical simulations of the dynamics. We employ a four-order Runge-Kutta integrator with time step $h/5$ and initial states sampled from a uniform distribution over the computational region $\Omega$. Then the dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ is constructed by taking snapshots of the generated trajectories at times $t_{j}$ and $t_{j}+h$, $1\leq j\leq M$. In total, there are $5\times 10^{5}$ points in the dataset for the first example ($4\times 10^{5}$ points for the second example). We build a representative subset $\{\tilde{X}_{k}\}$ to impose the orthogonality condition as in Eq. (7) using the Algorithm 1 of Ref. [21] with a parameter $r$ of $0.1$ for the first example and $5\times 10^{-4}$ for the second example.
@@ Appendix C Estimation of the normalization constant in the invariant distribution for Example 1
[…]
@@
where $I_{\alpha}(\cdot)$ denotes the modified Bessel function of the first kind [31].
Acknowledgments
The work of B. Lin was partially supported by A*STAR under its AME Programmatic programme: Explainable Physics-based AI for Engineering Modelling & Design (ePAI) [Award No. A20H5b0142]. PB acknowledges partial support from the European Union’s NextGenerationEU programme, in the framework of PRIN 2022, project DIMIN.
@@ Acknowledgments
Data supporting the findings in this manuscript are available upon request.

### v2

@@ 3.2 Example 2: Nanomechanical graphene resonator
Long-term prediction. Here we analyze the accuracy of the prediction for the identified dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$. Given $1000$ observed trajectories of the system (28) with initial states sampled from $\{\mathbf{x}\in\Omega:V_{\theta}(\mathbf{x})<\tau\}$, we generate trajectories by evolving the identified dynamics from the states. The statistics of the errors for the generated trajectories, as defined in Eq. (21), indicate that within one standard deviation, the error may be as high as approximately 5%. Fig. 9 (a) shows a comparison of the generated trajectories with the observed ones, demonstrating the capacity to accurately capture the underlying dynamical evolution.
Invariant distribution. With the identified quasipotential $U^{\text{Symb}}(P,Q)$ of Eq. (29), we compute an explicit invariant distribution for the nanomechanical resonator when perturbed by a random noise. It is important to investigate how fluctuations affect the out-of-equilibrium dynamics of nonlinear resonators because atomically thin resonators are very sensitive to environmental noise. The invariant distribution of the randomly perturbed system is approximated by
$p_{\epsilon}(P,Q)=Z_{\epsilon}^{-1}\exp(-U^{\text{Symb}}(P,Q)/\epsilon).$ (31)
Here the normalization constant is estimated by the integration over the domain $\Omega$. The evolution of the invariant distribution $p_{\epsilon}(P,Q)$ for different values of $\epsilon$, namely $10^{-9}$, $2\times 10^{-9}$ and $3\times 10^{-9}$, is showcased in Fig. 9 (b)-(d). While the predominance of the large oscillations of the nanomechanical resonator persists (motion around the large amplitude attractor - fixed point $A$), we observe a gradual smearing of the distribution in the phase space of the quadratures for larger $\epsilon$.
@@ 4 Conclusions
To conclude, we introduced a computational method for discovering explicit expressions of the quasipotential from observed data, which combines a neural network approximation and sparse regression. The main contribution of this paper is to demonstrate, for the first time, that it is possible to retrieve an analytical form of the quasipotential function using a fully data-driven approach. Two applications are presented. The first example validates the accuracy of the identified quasipotential, while the second highlights the ability of the data-driven method to interpret nanomechanical vibrational systems of practical engineering interest. Unlike existing approaches, our method can explicitly recover quasipotential expressions without any assumptions about the governing vector field. In particular, we do not use any a priori knowledge of the vector field equations. Instead, all information necessary to reconstruct the quasipotential function is derived directly from the data. This renders the method applicable to both synthetic and experimental datasets, as it only requires observations of the nonlinear dynamics. Moreover, the algorithm is robust to noisy data, which we demonstrate through tests mimicking realistic systems subject to perturbations and uncertainties.
The data-driven approach is suitable for a wide range of systems exhibiting complex multistable behavior, provided that sufficient trajectory data are available. Specifically, it requires adequate coverage of the phase space by the input trajectories to ensure accurate reconstruction. Additionally, it demands careful GPU memory management to avoid computational bottlenecks. Because our method relies on neural networks and sparse regression applied directly to available trajectory data rather than a discretized computational grid, its mathematical formulation and implementation fundamentally differ from grid-based methods. This avoids the need to operate on an n-dimensional grid, thereby mitigating the curse of dimensionality and facilitating extension to higher dimensions. We illustrate this capability with a four-dimensional example. Given these advantages, we believe the proposed approach offers a promising tool for precisely predicting key statistical quantities related to transitions between metastable states. This novel idea of determining a symbolic quasipotential function by coupling machine learning techniques presents a fully data-driven solution, demonstrating how recent advances in data science and machine learning open exciting new opportunities in the study of nonlinear dynamics.
Appendix A Initialization routine for the regularized regression
@@ Appendix B Data sampling and neural networks training
Sampling data
$\Omega$ $2,2]\!\times\![-1.5,1.5]\!\times\![-1.5,1.5] [$$0.02,0.02]\!\times\![-0.02,0.02]$
NN $5000$ $2000$
@@
$V_{\theta}^{\text{NN}}$
Structure 33-5050-5050-5050-11 22-100100-100100-100100-11
@@
$\mathbf{g}_{\theta}^{\text{NN}}$
Structure 33-5050-5050-5050-33 22-100100-100100-100100-22
@@
Training
Optimizer Adam Adam
@@
Data points are collected from $N$ trajectories generated from numerical simulations of the dynamics. We employ a four-order Runge-Kutta integrator with time step $h/5$ and initial states sampled from a uniform distribution over the computational region $\Omega$. Then the dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ is constructed by taking snapshots of the generated trajectories at times $t_{j}$ and $t_{j}+h$, $1\leq j\leq M$. In total, there are $5\times 10^{5}$ points in the dataset for the first example ($4\times 10^{5}$ points for the second example). We build a representative subset $\{\tilde{X}_{k}\}$ to impose the orthogonality condition as in Eq. (9) using the Algorithm 1 of Ref. [lin2021data] with a parameter $r$ of $0.1$ for the first example and $5\times 10^{-4}$ for the second example.
@@ Appendix C Estimation of the normalization constant in the invariant distribution for Example 1
$Z_{\epsilon}$ $=\iiint_{\mathbb{R}^{3}}\exp\big(-U^{\text{Symb}}(x,y,z)/\epsilon\big)dxdydz$ (36)
$=\int_{\mathbb{R}}\exp\left[-\frac{a_{1}}{\epsilon}x^{4}+\frac{a_{2}}{\epsilon}x^{2}-\frac{a_{3}}{\epsilon}\right]dx\cdot\int_{\mathbb{R}}\exp\left[-\frac{a_{4}}{\epsilon}y^{2}\right]dy\cdot\int_{\mathbb{R}}\exp\left[-\frac{a_{5}}{\epsilon}z^{2}\right]dz$
$=\int_{\mathbb{R}}\exp\left[-\frac{a_{1}}{\epsilon}x^{4}+\frac{a_{2}}{\epsilon}x^{2}-\frac{a_{3}}{\epsilon}\right]dx\cdot\sqrt{\pi\epsilon/a_{4}}\cdot\sqrt{\pi\epsilon/a_{5}}$
@@
where $I_{\alpha}(\cdot)$ denotes the modified Bessel function of the first kind [abramowitz1968handbook].
Appendix D Statistics of the learned coefficients for Example 2
Fig. 10 showcases the distributions of the learned coefficients for each library term in $U^{\text{Symb}}(\mathbf{x})$.
Figure 10: Distributions of the coefficients for each library term in the identified potential $U^{\text{Symb}}(P,Q)$ in $50$ independent runs for of Sec. 3.2.
Appendix E Error estimation for Example 2
In Fig. 11 we report the error between the reference quasipotential estimated using the standard ordered upwind method (Fig. 8 (c)) and the two data-driven steps of our method, namely the black-box neural network function, and the analytical expression obtained via symbolic regression.
Figure 11: Error in the estimated data-driven quasipotential. (a) Error of the neural network estimate (Fig. 8 (d)) relative to the ordered upwind method (Fig. 8 (c)). (b) Error of the symbolic regression estimate (Fig. 8 (e)) relative to the ordered upwind method (Fig. 8 (c)).
As shown in Fig. 11, the largest discrepancies occur in the outer regions of the phase space, where the decomposition is challenging to estimate due to the high flow speeds of the underlying vector field.
Acknowledgments and Funding Declaration
The work of BL was partially supported by A*STAR under its AME Programmatic programme: Explainable Physics-based AI for Engineering Modelling & Design (ePAI) [Award No. A20H5b0142]. PB acknowledges partial support from the European Union’s NextGenerationEU programme, in the framework of PRIN 2022, project DIMIN.
@@ Appendix E Error estimation for Example 2
All computational results are reproducible and code can be found at https://github.com/LinBoNUS/SIQ.
@@
Author Contributions
P.B. and B.L. conceived the research idea. B.L. developed the code and conducted the numerical investigations. P.B. supervised the project. All authors contributed to data analysis, interpretation of the results, and writing of the manuscript.